we want to mix so many pounds of almonds at $4.27 / pound
and so many pounds of cashews at $3.10 / pound
to get our final mix consisting of 60 pounds @ $3.49 / pound
how many pounds of almonds and cashews do i need
i'm having problems setting this up...here's what i have so far...
Type Cost per pound pounds Total cost
Almonds $4.27 x $4.27(x) =?
Cashews $3.10 y $3.10(y)=?
Mixture $3.49 60=x+y 3.49(60) = 209.40
so...i need to substitue x or y ?
so i take 60=x+y and solve for x
x=y-60
now i take my x term and substitute 4.27(y-60) = ?
ok...i need some help. Thx

Answer by CharStar(110)

You are close, but you don't need a y. Just an x. See if the following will be some help to you.
Almonds - $4.27
Amount in lbs = x
Total of Each - 4.27x
Cashews - $3.10
Amount in lbs = 60-x
Total Cost of each = $3.10(60-x)
Mixture =$3.49
Amount in lbs = 60
Total Cost of each = 3.49(60)
4.27x + 3.10(60-x) = 3.49(60)
4.27x + 186 - 3.10x = 209.40
1.17x = 23.40
x=20
Check your answer
4.27(20)85.40
$3.10(60-20)$3.10 x 40 = 124.00
85.40 + 124.00 = 209.40